the weight of a cylindrical metal varies jointly as its length Land the square of the diameter d of its base if w=6kg when L=6cm and d=3cm find the equation of the variation.
Answer by stanbon(75887)

the weight of a cylindrical metal varies jointly as its length Land the square of the diameter d of its base if w=6kg when L=6cm and d=3cm find the equation of the variation.
-----------
w = k*L*d^2
-----
6 = k*6*3^2
6 = 54k
k = 1/9
------
Equation of variation:
w = (1/9)*L*d^2
-------------------
Cheers,
Stan H.
-------------------